# Neural Network-Based Bandit: A Medium Access Control for the IIoT Alarm Scenario

### v1

@@ 15.  Transmission pattern $\boldsymbol{a}_{\upsilon}$, $\varepsilon$, $\mathbf{w}$
III-F MAC IN NNBB
The MAC of an IIoT device in NNBB is described by three states: Normal State (NS), Emergency State (ES), and Quiet State (QS). In NS, the IIoT device transmits data from the sensed process. In ES, an active device transmits the alarm message. In QS, an inactive device halts its process data transmission. The latter can be resumed in a future time slot, provided the device is no longer in QS.
@@ III-F MAC IN NNBB
Note that our target is to improve the success rate of the first transmission attempt of the alarm message, rather than focusing on its retransmissions. Thus, we skipped the alarm retransmission procedure.
@@
IV NUMERICAL RESULTS
We now assess the proposed NNBB RA scheme through simulations. We consider a circular region around the BS, where the ExC and the devices are uniformly distributed. The circular region has a fixed device density of $0.2$ devices///m2, and we vary the radius of the area to accommodate the specified number of devices. Similar to [4], we use
@@ IV NUMERICAL RESULTS
as the activation probability function, where $\lambda$ is a mean-scaling multiplier. The channels undergo Rayleigh fading and ${\mathbf{c}_{\upsilon}\sim\mathcal{CN}(\mathbf{0},\frac{1}{r_{\upsilon}^{\gamma}}\mathbf{I}_{M})}$, where ${r_{\upsilon}^{\gamma}}$ represents the path loss, $\gamma$ is the path loss exponent, and $r_{\upsilon}$ is the distance between the device $\upsilon$ and the BS. The simulation parameters and their corresponding values are listed in Table III. We measure the performance of a RA scheme using the success rate metric, which is defined as the probability that an alarm event is successfully reported to the ExC. We measure the success rate of an algorithm after it has converged. To illustrate the variability in the success rate of the first transmission attempt in our simulation results, we generate an alarm at each time slot and discard the respective alarm message from the active devices after its first transmission attempt.
IV-A BENCHMARK SCHEMES
We use MAB-based RA [31] and RS as benchmarks. Akin to NNBB, MAB and RS are implemented at each IIoT device. MAB uses the $\varepsilon$-greedy method to select an action when an alarm event is detected. The action values, $Q(\boldsymbol{\mathring{a}}_{i})\ \forall i=1,\cdots,2^{M}$, in the MAB-based RA scheme are updated as
$Q(\boldsymbol{\mathring{a}}_{i})\leftarrow(1-\tau)Q(\boldsymbol{\mathring{a}}_{i})+r(\boldsymbol{\mathring{a}}_{i})\tau,$ (15)
where $\tau$ is the learning rate with an initial value of $1$ and it decreases with time. In contrast with [6], we simultaneously train the MABs of all the active devices at each alarm event, i.e., the action values of every active device are updated. On the other hand, in the RS scheme, actions are selected randomly at each alarm event.
@@ IV-A BENCHMARK SCHEMES
IV-B IMPACT OF THE NUMBER OF DEVICES AND CHANNELS
Fig. 5 shows the impact of the number of devices on the success rate achieved by NNBB and the benchmark schemes. As the number of devices increases, the success rate deteriorates. However, NNBB increasingly outperforms MAB and RS. Specifically, for a network of $20$ devices with $5$ channels, the success rate of NNBB is $9\%$ higher than that of MAB. This gain increases to $20\%$ when there are $60$ devices. Although MAB outperforms NNBB when only $2$ channels are shared by at most $20$ devices, the superiority is marginal. The superiority disappears and gets reversed as the number of devices increases. All in all, as the number of devices increases, the performance gain of NNBB over MAB and RS increases.
Fig. 6 plots the success rate as a function of the number of channels. We observe that NNBB outperforms MAB and RS regardless of the number of available channels for heavily loaded networks. Notably, increasing the number of available channels from $2$ to $6$ results in a $30\%$ increase in the success rate for NNBB in a network of $60$ devices, whereas MAB only experiences a $13\%$ improvement. A closer examination of the plots in Fig. 6 also reveals that NNBB is less sensitive to an increase in the number of devices than MAB and RS.
@@ IV-B IMPACT OF THE NUMBER OF DEVICES AND CHANNELS
IV-C IMPACT OF THE ACTIVATION PROBABILITY
Fig. 7 illustrates the effect of $\lambda$, which directly impacts the activation probability, on the performance of the RA schemes. As $\lambda$ increases, the activation probability increases. The figure indicates that NNBB and the benchmark schemes exhibit similar success rates when the activation probability is small, i.e., when $\lambda=1$. However, as the activation probability increases, the success rate of NNBB decays at a slower rate compared to MAB and RS. Even for $\lambda=3$, when the success rates of the benchmark schemes fall well below $0.8$, the success rate of NNBB remains significantly higher. Notably, for NNBB, a significant difference in the success rates between the two considered networks is visible only when $\lambda=4$, while for MAB and RS the difference becomes apparent already when $\lambda=3$.
IV-D IMPACT OF THE NUMBER AND SIZE OF THE HIDDEN LAYERS
@@
Fig. 8 illustrates the performance of NNBB as a function of the number of hidden layers and their size. It is evident from the figure that NNBB consistently outperforms the benchmark approaches. Interestingly, the hidden layer combination with the least number of neurons, i.e., ${(H,h)=(2,1)}$, yields the highest success rate among the considered configurations. Additionally, this configuration outperforms the other two configurations throughout the training procedure, as indicated in Fig. 9. It is often advantageous to reduce the hidden layer size as much as possible to maintain the NN’s generalization capability. The excess neurons can cause the network to act like a memory bank, leading to suboptimal performance when presented with inputs other than the training samples [32]. This justifies the suboptimal performance of the configurations ${(H,h)=(1,10)}$ and ${(H,h)=(2,15)}$. Reducing the hidden layer size also reduces the computational complexity of NNBB. Furthermore, Fig. 9 demonstrates that for a network of $40$ devices, NNBB should train over $15200$ and $31000$ alarm events to converge when $5$ and $6$ channels, respectively, are available.
@@
V CONCLUSION
We proposed NNBB, which is a distributed DRL-based RA scheme that allows IIoT devices to develop implicit coordination to successfully convey an alarm message to an external controller. Specifically, upon the detection of an alarm event, every active device starts a procedure to acquire a useful context to feed a local DNN. Then, with the help of the DNN and the $\varepsilon$-greedy method, a device selects the transmit channel(s) for the alarm message, including also the possibility of no transmission. A reward or penalty is granted based on the success or failure of the transmission, which eventually is used for training the DNN. Notably, the DNN uses two hidden layers with just one neuron each, significantly decreasing NNBB’s computational complexity. Simulations show that NNBB experiences a relatively lower drop in its success rate compared to the benchmark schemes as the number of devices in a network increases. Further, the success rate gained by increasing the available channels is relatively higher.
For future work, it might be interesting substituting the adopted $\varepsilon$-greedy method, which is less effective in large action spaces and sensitive to the initial value of $\varepsilon$, by 1) Thompson sampling-based exploration, and 2) upper confidence bound-based exploration. Meanwhile, NNBB with the help of context develops implicit coordination among devices. Therefore, exploiting NNBB in connected robotics and autonomous systems are interesting future research directions.

### v2

@@ 15.  Transmission pattern $\boldsymbol{a}_{\upsilon}$, $\varepsilon$, $\mathbf{w}$
Note that randomly sampling a mini-batch from the memory buffer breaks the temporal correlation between samples collected sequentially. Additionally, updating the DNN using mini-batches leads to DNN updates that are averaged over multiple samples. This reduces the variance of DNN updates across time compared to the case where the DNN is updated with a single sample, making the training process of DNN stable. Furthermore, using mini-batches helps prevent agents from overfitting to short-term fluctuations. Notably, it is important to choose an appropriate mini-batch size. A too-small mini-batch size may result in noisy DNN updates, causing oscillations in the RA policies of agents. On the other hand, a too-large mini-batch size can make DNN updates computationally expensive and may also increase the chances of overfitting to the sampled mini-batch.
III-E ALGORITHM OUTLINE AND MAC ASPECTS
  1. 1.
    Transmit their respective pilot signals to the BS, which then broadcasts the received aggregated signal back to them. With this, each active agent acquires its context.
  2. 2.
    Feed their context to their DNN, which yields $2^{M}$ action values.
  3. 3.
  4. 4.
    Monitor the feedback from the ExC and determine its reward.
  5. 5.
  6. 6.
    Train their DNNs.
Algorithm 1 provides a comprehensive description of NNBB.
The IIoT device operation comprises three states: Normal State (NS), Emergency State (ES), and Quiet State (QS)[35]. In NS, the IIoT device transmits data from the sensed industrial process to the BS. In ES, an active device transmits the alarm message to the ExC. In QS, an inactive device halts its process data transmission to the BS. The latter can be resumed in a future time slot, provided the device is no longer in QS. Note that the primary duty of the BS is to collect process data from IIoT devices, while its secondary duty is to broadcast the context.
@@ III-E ALGORITHM OUTLINE AND MAC ASPECTS
Note that our target is to improve the success rate of the first transmission attempt of the alarm message, rather than focusing on its retransmissions. Thus, we have skipped the alarm retransmission procedure.
IV THEORETICAL AND PRACTICAL INSIGHTS
This section offers comprehensive discussions on the complexity, convergence, and robustness to system dynamics of the proposed NNBB.
IV-A COMPUTATIONAL COMPLEXITY
We quantify the computational complexity of NNBB in terms of the number of arithmetic operations executed for making a single alarm transmission decision. The complexities of fundamental operations, such as ReLU and operations in steps 234, and 6 of Algorithm 1, are $1,1,1,{|\mathbf{\mathring{A}}|}$, and $1$, respectively, where ${|\mathbf{\mathring{A}}|=2^{M}}$.
The complexity of NNBB accounts for the complexities across three distinct phases: action values generation phase, action selection phase, and training phase. The complexity for the first and third phase is ${\vartheta_{1}=\sum_{i=1}^{|\boldsymbol{l}|-1}l_{i+1}(2l_{i}+1)}$ and ${\vartheta_{3}=B\vartheta_{1}}$, respectively, as derived in [36]. Here, ${\boldsymbol{l}=[l_{1},\cdots,l_{|\boldsymbol{l}|}]^{T}}$ with ${l_{1}=|\boldsymbol{s}_{\upsilon}|}$ and ${l_{|\boldsymbol{l}|}=|\mathbf{\mathring{A}}|}$, while the remaining elements of ${\boldsymbol{l}}$ are the hidden layer sizes. Moreover, because of steps 36 of Algorithm 1, the complexity of the second phase falls within the range ${[3,2+|\mathbf{\mathring{A}}|]}$. According to Table III, we have ${\boldsymbol{l}=[M,1,1,2^{M}]}$. Thus, the lower bound complexity $(\vartheta_{lb})$ and upper bound complexity $(\vartheta_{ub})$ expressions for NNBB are
$\vartheta_{lb}=$ $\ \vartheta_{1}+\vartheta_{3}+3$
$=$ $\ (B+1)\vartheta_{1}+3$
$=$ $\ 90\times 4^{M}+(123+60M)2^{M}+2M+7,$ (14a)
$\vartheta_{ub}=$ $\ \vartheta_{1}+\vartheta_{3}+2+|\mathbf{\mathring{A}}|$
$=$ $\ \vartheta_{lb}+2^{M}-1.$ (14b)
[…]
$\vartheta$ $=O(4^{M}).$ (15)
[…]
Consider an IIoT network with a large ${|\mathcal{N}|}$, thus, potentially large ${|\mathcal{N}^{\prime}|}$ depending on the device density. A large ${|\mathcal{N}^{\prime}|}$ increases the likelihood of collisions when $M$ is small, minimizing the chances of a successful alarm transmission. Meanwhile, a large $M$ can increase the chances of a successful alarm transmission. However, it would also exponentially increase the complexity of NNBB, hindering NNBB’s implementation on resource-scarce IIoT devices. An appealing solution to these problems lies in: (i) dividing the large IIoT network into several small sub-networks, (ii) partitioning the large set of orthogonal channels into several smaller subsets, and (iii) allocating a subset to each of these sub-networks, ensuring that the value of $M$ within each sub-network remains small. The network division approach from [4], where IIoT devices with high spatial correlation are kept in different sub-networks, can be an effective method for forming sub-networks. This way (i) NNBB implemented on an IIoT device would only take into account the value of $M$ assigned to the sub-network in which the respective IIoT device belongs, and (ii) the chances of a successful alarm transmission would certainly increase. By following this implementation approach, NNBB can efficiently manage alarm transmissions in large IIoT networks, without increasing its complexity.
IV-B CONVERGENCE
In a multi-agent distributed RL system, where multiple agents independently train their DNNs towards a common goal, ensuring convergence becomes a complex challenge. An action-selection strategy can significantly influence the stability and convergence of such a system by ensuring a balance between exploration and exploitation [6]. In this work, we are employing one of the most commonly used action-selection strategies, which is $\varepsilon$-greedy. By carefully scheduling the decrease of $\varepsilon$, while ensuring (i) it remains high for long enough to adequately explore the action space and (ii) its minimum value is high enough to allow some level of exploration, we can prevent abrupt shifts in agents’ policies, which could otherwise lead to instability. Thus, a well-tuned $\varepsilon$-greedy strategy promotes smooth convergence in the multi-agent distributed RL system. To satisfy aforesaid conditions, we chose to gradually decrease $\varepsilon$ from $1$ to $0.1$ with a step size of $0.005$ after every alarm event.
Meanwhile, gradient clipping plays an essential role in stabilizing the training process of the DNN, which in turn can facilitate the convergence of our multi-agent distributed RL system. By limiting the maximum value of gradients during backpropagation, gradient clipping helps maintain a balanced learning process across all agents, preventing any single agent’s updates from becoming disproportionately large, which could disrupt the collective learning process. Additionally, large gradients can exacerbate the sensitivity to the learning rate of the DNN, potentially leading to instability if the learning rate is not finely tuned. Gradient clipping reduces this sensitivity, making the training process of the DNN more robust to different learning rate settings. Furthermore, gradient clipping accelerates the NN training [37]. The effectiveness of gradient clipping depends on the global norm (of the gradient vector) threshold $\beta_{0}$. If $\beta_{0}$ is too high, gradient clipping will rarely activate, causing the RA policies of agents to oscillate and struggle to converge. Conversely, if $\beta_{0}$ is too low, agents may stuck with suboptimal RA policies, as the small gradient updates could stop agents from exploring the policy space. To leverage the aforesaid benefit provided by gradient clipping, we chose to implement it in NNBB with ${\beta_{0}=5}$.
The mean square error of the system, at an alarm instance, is defined as
$\textrm{MSE}_{\textrm{sys}}=\frac{1}{|\mathcal{N}^{\prime}|}\sum_{\upsilon\in\mathcal{N}^{\prime}}\Omega_{\upsilon}.$ (16)
Fig. 5 illustrates the variation in the mean square error of the multi-agent system during the training phase of NNBB. As can be seen in Fig. 5, the ${\textrm{MSE}_{\textrm{sys}}}$ is decreasing consistently with the progression of NNBB’s training across agents, regardless of the DNN configuration. Furthermore, ${\textrm{MSE}_{\textrm{sys}}}$ is becoming stable at the end for all the DNN configurations, signifying the convergence of our multi-agent distributed RL system. Thus, Fig. 5 proves that utilizing the right action-selection strategy and gradient clipping can facilitate the convergence of NNBB, irrespective of the system configuration.
Figure 5: Variation in ${\textrm{MSE}_{\textrm{sys}}}$ during the training phase of NNBB for ${|\mathcal{N}|=20}$ and ${M=4}$.
IV-C ROBUSTNESS TO SYSTEM DYNAMICS
[…]
Moreover, note that each agent trains its own DNN model independently in NNBB. The DNN is designed to handle complex and time-varying inputs, meaning it could learn to recognize patterns in the input data that correlate with channel fading. This allows agents to learn and adapt to time-varying channel fading based on their local experiences, thus, making NNBB resilient to variations in channel fading. We have assumed Rayleigh fading for performance analysis in Section VI, which is one of the worst types of fading due to a lack of a deterministic line-of-sight component. Therefore, in the case of less severe (less random) fading environments, such as those with Rician or Nakagami-m $(m>1)$ fading, the performance of NNBB would certainly be better than the one obtained with Rayleigh fading.
V BENCHMARK SCHEMES
We use MAB-based RA [38], RS, and MQLFA-based RA as benchmarks. Akin to NNBB, MAB, RS, and MQLFA are implemented at each IIoT device. In the RS scheme, actions are selected randomly at each alarm event, constituting a typical dumb approach. Next, we comprehensively discuss the more intelligent MAB and MQLFA benchmark schemes.
V-A MAB-BASED RA
MAB uses the $\varepsilon$-greedy method to select an action when an alarm event is detected. Note that the update strategy for $\varepsilon$ in MAB-based RA is the same as the one adopted for NNBB. The action values, ${Q_{1}(\boldsymbol{\mathring{a}}_{i})}$, ${\forall i=1,\cdots,2^{M}}$, in the MAB-based RA scheme are updated as
$Q_{1}(\boldsymbol{\mathring{a}}_{i})\leftarrow(1-\tau)Q_{1}(\boldsymbol{\mathring{a}}_{i})+r(\boldsymbol{\mathring{a}}_{i})\tau.$ (17)
Here, $\tau$ is the learning rate. It has an initial value of $1$ and decreases with time. In contrast with [6], we simultaneously train the MABs of all the active devices at each alarm event, i.e., the action values of every active device are updated.
V-B MQLFA-BASED RA
MQLFA is an extension of the traditional Q-learning algorithm where each agent learns its own RA policy independently and uses the linear function approximation to compute action values as well as to handle the continuous state space. Specifically, the action value is represented as a linear combination of features that describe the state-action pair, i.e., a weighted sum of feature values. Thus, action values, ${Q_{2}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i},\boldsymbol{\theta}_{\upsilon})}$, ${\forall i=1,\cdots,2^{M}}$, in MQLFA are computed as
$Q_{2}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i},\boldsymbol{\theta}_{\upsilon})=\boldsymbol{\theta}_{\upsilon}^{T}\boldsymbol{\phi}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i}),$ (18)
where ${\boldsymbol{\phi}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i})=[\operatorname{diag}(\tilde{\boldsymbol{s}}_{\upsilon}\tilde{\boldsymbol{s}}_{\upsilon}^{H}),\ \boldsymbol{\mathring{a}}_{i}]^{T}\in\mathbb{R}^{2M\times 1}}$ is a feature vector, $\tilde{\boldsymbol{s}}_{\upsilon}$ is the normalized $\boldsymbol{s}_{\upsilon}$ and is computed as
$\tilde{\boldsymbol{s}}_{\upsilon}=\frac{\overbrace{\boldsymbol{s}_{\upsilon}-\min(\textrm{Re}(\boldsymbol{s}_{\upsilon}))-\min(\textrm{Im}(\boldsymbol{s}_{\upsilon}))j}^{\boldsymbol{\kappa}}}{\max(\textrm{abs}(\boldsymbol{\kappa}))}.$ (19)
Moreover, ${\boldsymbol{\theta}_{\upsilon}\in\mathbb{R}^{2M\times 1}}$ is a weight vector and each agent learns its ${\boldsymbol{\theta}_{\upsilon}}$ independently.
In MQLFA-based RA, after detecting an alarm, active agents first transmit their respective pilot signals and receive the context $\boldsymbol{s}_{\upsilon}$. Then, active agents compute their respective action values, ${Q_{2}(\boldsymbol{s}_{\upsilon},\boldsymbol{\mathring{a}}_{i},\boldsymbol{\theta}_{\upsilon})}$, ${\forall i=1,\cdots,2^{M}}$, ${\forall\upsilon\in\mathcal{N}^{\prime}}$, and utilize them to select their respective actions, $\boldsymbol{a}_{\upsilon}$, ${\forall\upsilon\in\mathcal{N}^{\prime}}$, with the help of the $\varepsilon$-greedy method. Note that the update strategy for $\varepsilon$ in MQLFA-based RA is the same as the one adopted for NNBB. Next, after receiving the reward ${r(\boldsymbol{a}_{\upsilon})}$, active agents update their respective ${\boldsymbol{\theta}_{\upsilon}}$, ${\forall\upsilon\in\mathcal{N}^{\prime}}$, as
$\boldsymbol{\theta}_{\upsilon}\leftarrow\boldsymbol{\theta}_{\upsilon}+(r(\boldsymbol{a}_{\upsilon})-Q_{2}(\boldsymbol{s}_{\upsilon},\boldsymbol{a}_{\upsilon},\boldsymbol{\theta}_{\upsilon}))\tau\boldsymbol{\phi}(\boldsymbol{s}_{\upsilon},\boldsymbol{a}_{\upsilon}).$ (20)
Note that MQLFA-based RA takes into account $\boldsymbol{s}_{\upsilon}$. This makes the MQLFA-based RA scheme a suitable benchmark for our proposed NNBB scheme.
@@ V-B MQLFA-BASED RA
Initial value of ${\boldsymbol{\theta}_{\upsilon}}$ $\mathbf{0}_{2M}$
VI NUMERICAL RESULTS
We now assess the performance of the proposed NNBB RA scheme through simulations. We consider a circular region around the BS, where the ExC and devices are uniformly distributed. The circular region has a fixed device density of $0.2$ devices///m2, and we vary the radius of the area to accommodate the specified number of devices. Similar to [4], we use
@@ VI NUMERICAL RESULTS
as the activation probability function, where ${\lambda>0}$ is a mean-scaling multiplier.
Pilot symbols ${\{\varrho_{i,\upsilon}|\forall i=1,\cdots,M\}}$ are sampled from the symmetric quadrature phase shift keying constellation; i.e., chosen uniformly at random from =${\mathcal{B}=\{\frac{1+j}{\sqrt{2}},\frac{-1-j}{\sqrt{2}},\frac{1-j}{\sqrt{2}},\frac{-1+j}{\sqrt{2}}\}}$. Note that pilot symbols ${\{\varrho_{i,\upsilon}|\forall i=1,\cdots,M\}}$ are independent and identically distributed random variables with probability mass function ${\{\textrm{Pr}(\varrho_{i,\upsilon}=x)=\frac{1}{4}|x\in\mathcal{B},\forall i=1,\cdots,M\}}$. The channel undergoes quasi-static Rayleigh fading and ${\mathbf{c}_{\upsilon}\sim\mathcal{CN}(\mathbf{0},\frac{1}{r_{\upsilon}^{\gamma}}\mathbf{I}_{M})}$, where ${r_{\upsilon}^{\gamma}}$ represents the path loss, $\gamma$ is the path loss exponent, and $r_{\upsilon}$ is the distance between the device $\upsilon$ and the BS. The simulation parameters and their corresponding values are listed in Table III.
We measure the performance of an RA scheme using the success rate metric, which is defined as the probability that an alarm event is successfully reported to the ExC. We measure the success rate of an algorithm after it has converged. To illustrate the variability in the success rate of the first transmission attempt in our simulation results, we generate an alarm at each time slot and discard the respective alarm message from the active devices after its first transmission attempt.
@@
VI-A IMPACT OF THE NUMBER OF DEVICES AND CHANNELS
Fig. 6 shows the impact of the number of devices on the success rate achieved by NNBB and benchmark schemes. As the number of devices increases, the success rate deteriorates. However, NNBB increasingly outperforms MAB, RS, and MQLFA. Specifically, for a network of $20$ devices with $5$ channels, the success rate of NNBB is $9\%$ higher than that of MAB. This gain increases to $20\%$ when there are $60$ devices. Although MAB outperforms NNBB when only $2$ channels are shared by at most $20$ devices, the superiority is marginal. The superiority disappears and gets reversed as the number of devices increases. Meanwhile, NNBB outperforms MQLFA for all network configurations. Even though both NNBB and MQLFA utilize context, NNBB processes it more effectively due to the use of the DNN. This allows NNBB to compute relatively better action values than MQLFA, which later leads to relatively better action selection. Moreover, the utilization of context in NNBB and MQLFA makes them more robust to an increase in the number of devices compared to MAB and RS. This allows MQLFA to surpass MAB in terms of success rate when the number of devices goes beyond $40$ in some network configurations. All in all, as the number of devices increases, the performance gain of NNBB over MAB, RS, and MQLFA increases.
Fig. 7 plots the success rate as a function of the number of channels. We observe that NNBB outperforms MAB, RS, and MQLFA regardless of the number of available channels for heavily loaded networks. Notably, increasing the number of available channels from $2$ to $6$ results in a $30\%$ increase in the success rate for NNBB in a network of $60$ devices, whereas MAB only experiences a $13\%$ improvement. A closer examination of the plots in Fig. 7 also reveals that NNBB is less sensitive to an increase in the number of devices than MAB, RS, and MQLFA. Moreover, the exploitation of context in NNBB and MQLFA is making them more robust than MAB and RS to an increase in the number of channels.
@@ VI-A IMPACT OF THE NUMBER OF DEVICES AND CHANNELS
VI-B IMPACT OF THE ACTIVATION PROBABILITY
Fig. 8 illustrates the effect of $\lambda$, which directly impacts the activation probability, on the performance of the RA schemes. As $\lambda$ increases, the activation probability increases. The figure indicates that NNBB and benchmark schemes exhibit similar success rates when the activation probability is small, i.e., when $\lambda=1$. However, as the activation probability increases, the success rate of NNBB decays at a slower rate compared to MAB, RS, and MQLFA. Even for $\lambda=3$, when the success rates of benchmark schemes fall well below $0.8$, the success rate of NNBB remains significantly higher. Notably, for NNBB, a significant difference in the success rates between the two considered networks is visible only when $\lambda=4$, while for benchmark schemes the difference becomes apparent already when $\lambda=3$.
VI-C IMPACT OF THE NUMBER AND SIZE OF THE HIDDEN LAYERS
@@
Fig. 9 illustrates the performance of NNBB as a function of the number of hidden layers and their size. It is evident from the figure that NNBB consistently outperforms the benchmark approaches. Interestingly, the hidden layer combination with the least number of neurons, i.e., ${(H,h)=(2,1)}$, yields the highest success rate among the considered configurations. Additionally, this configuration outperforms the other two throughout the training procedure, as indicated in Fig. 10. It is often advantageous to reduce the hidden layer size as much as possible to maintain the NN’s generalization capability. Excess neurons can cause the network to act like a memory bank, leading to suboptimal performance when presented with input other than training samples [40]. This justifies the suboptimal performance of the configurations ${(H,h)=(1,10)}$ and ${(H,h)=(2,15)}$. Reducing the hidden layer size also reduces the computational complexity of NNBB. Furthermore, Fig. 10 demonstrates that for a network of $40$ devices, NNBB should train over $15200$ and $31000$ alarm events to converge when $5$ and $6$ channels, respectively, are available.
@@
VII CONCLUSION
We proposed NNBB, a distributed DRL-based RA scheme that allows IIoT devices to develop implicit coordination to successfully convey an alarm message to an external controller. Specifically, in the proposed scheme, upon the detection of an alarm event, every active device starts a procedure to acquire a useful context to feed a local DNN. Then, with the help of the DNN and the $\varepsilon$-greedy method, the device selects the transmit channel(s) for the alarm message, including also the possibility of no transmission. A reward or penalty is granted based on the success or failure of the transmission, which eventually is used to train the DNN. Notably, the proposed DNN employs two hidden layers with just one neuron each, significantly decreasing NNBB’s computational complexity. Simulations revealed that NNBB experiences a relatively lower drop in its success rate compared to the benchmark schemes as the number of devices in a network increases, while the success rate gained by increasing the available channels is relatively higher.
For future work, it might be interesting to substitute the adopted $\varepsilon$-greedy method, which is less effective in large action spaces and sensitive to the initial value of $\varepsilon$, by Thompson sampling-based exploration or upper confidence bound-based exploration. Meanwhile, NNBB with the help of context develops implicit coordination among devices. Therefore, exploiting NNBB in connected robotics and autonomous systems is an interesting future research direction.